OpenStudy (anonymous):

Use the given information to evaluate each expression. cos(φ) = 4/5,(0° < φ < 90°)

OpenStudy (anonymous):

do you have any idea how to start?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

what are we supposed to be looking for here?

OpenStudy (anonymous):

are we looking for the angle, the other trig functions, what??

OpenStudy (anonymous):

sorry sin(2φ)

OpenStudy (anonymous):

cos(2φ) tan(2φ)

OpenStudy (anonymous):

i got it give me a sec

OpenStudy (anonymous):

so i got: \[\sin2\phi=24/25\] \[\cos2\phi=7/25\] \[\tan2\phi=24/7\]

OpenStudy (anonymous):

and they are all positive since the values lie in the first quadrent

OpenStudy (anonymous):

can you should me how you got the answer?

OpenStudy (anonymous):

okay first we have to extablis something, we are looking for \[0\le2\phi \le90\]. Solving this inequality we get:\[0\le \phi \le45\]

OpenStudy (anonymous):

This means that we are working in the first quadrent, which denotes that all trig functions are positive okay

OpenStudy (anonymous):

Now, we want sin2x, cos2x, and tan2x. These allow us to use the double angle fromulas. sin2x=2sinxcosx cos2x=cos^2x-sin^2x and Tan2x=(2tanx)/(1-tan^2x)

OpenStudy (anonymous):

now we also know that we are given taht cosx=4/5, which should automatically denote a 3-4-5 triangle. This will then make sinx=3/5 and tanx=3/4

OpenStudy (anonymous):

Now all we havet to do is crunch out the equations plugging in the corresponding values into each double angle formula

OpenStudy (anonymous):

sin2x=2(3/5)(4/5)=24/5

OpenStudy (anonymous):

cos2x=(4/5)^2-(3/5)^2=16/25-9/25=7/25

OpenStudy (anonymous):

tan2x=2(3/4)/2-(3/4)^2=(3/2)/(1-9/16)=(3/2)*(16/7)=24/7

OpenStudy (anonymous):

and thats that

OpenStudy (anonymous):

satisfied??

OpenStudy (anonymous):

yes. Thank you.